# Dominik Bojko Krzysztof Grining Marek Klonowski

### v6

@@ 3 Differential Privacy Preliminaries
In this section, we briefly recall differential privacy. For more details, see, e.g. [18]. We denote the set of (positive) natural numbers by $\mathbb{N}$ and the set of all integers by $\mathbb{Z}$. Moreover, let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $a,b\in\mathbb{Z}$ let us define a discrete interval $a,b]\cap\mathbb{Z}$ by $[a:b]$. We also define $n]=\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$. We assume that there exists a trusted curator who holds, or securely obtains, the data of individuals in a (possibly distributed) database $x$. Every row of $x$ consists of the data of some individual. By $\mathcal{X}$, we denote the space of all possible rows. The goal is to protect the data of every single individual, even if all users except one collude with an adversary to breach the privacy of this single, uncorrupted user. On the other hand, the curator is responsible for producing a release – a possibly accurate response to a requested query. This response is then released to the public, who is allowed to perform a statistical analysis on it. The differential privacy is, by design, resilient to post-processing attacks, so even if the adversary obtains the public release, he will not be able to infer anything about specific individuals participating in this release.
@@
Definition 2 (Differential Privacy (formulation from [18]))
@@
Then $\mathbb{E}\left(X\right)=0$ and $\mathrm{Var}(X)=2\lambda^{2}$.
Let $c(x)$ be the number of rows in $x$, which satisfy a given property. Note that $c$ in the differential privacy literature is usually referred to as count query. Imagine that an aggregating mechanism is defined as follows: $\mathcal{M}(x)=c(x)+\mathcal{L}(\varepsilon^{-1})$. Then $\mathcal{M}$ is $(\varepsilon)$-DP (for more precise properties of Laplace noise, see [18]).
In the privacy analysis of large-scale distributed protocols, two types of approaches are typically distinguished: event/record-level (e.g., [51, 42]), where information about a specific event/record is protected, and user-level privacy (e.g., [34]), which protects the privacy of the users themselves. Generally, the latter is a stronger model (see the discussion in [34]).
@@ 4 Probabilistic Counters — preliminaries
In Figure 1, one can see a graphical representation of the probabilistic counter. As mentioned, increment requests are indicated by $^{\prime}1^{\prime}$ and other rows by $^{\prime}0^{\prime}$ input. The dice represent the randomness. The $\mathbf{X}$-mark indicates that there is no action.
@@ 4.1 Morris Counter
We begin with a short description of the Morris Counter (originally referred to as an approximate counter [36, 24]). Fix $a>1$. The algorithm 1 is a very simple pseudocode of the Morris Counter [36].
@@ 5    $M\leftarrow M+1$;
Roughly speaking, we start with $M=1$. Each incoming increment request triggers a random event. This event increments the counter ($M\leftarrow M+1$) with probability $a^{-M}$ ($r\sim\mathrm{Uni}([0,1])$ generates a number uniformly at random from the interval $[0,1]$, using in practice some Pseudo Random Number Generator). Note that this approximate counting protocol can be easily distributed. Indeed, any entity who wants to increment the counter only has to send the request to increment it. These requests can be queued on the server and resolved one after another. A detailed description of the approximate counting method can be found in [36, 24]. Throughout this article, we examine only a standard Morris Counter i.e., with the base $a=2$. Morris Counter can also be defined recursively.
@@
The accuracy of Morris Counter has been thoroughly analysed in various classical papers. The first detailed analysis was proposed by Ph. Flajolet in [24]. In this part, we present the essence of theorems presented in this paper, which will be useful later on.
@@ 4  $C\leftarrow\max\{C,r\}$;
The expectation and variance of the maximum of $n$ i.i.d. geometric variables have already been analysed in the literature. For instance, Szpankowski and Rego [47] provided exact formulas for the’ expected value and variance of such variables. However, they are impractical for large applications $n$. Hence, they also provided asymptotics (here, for a maximum of $n$ independent $\mathrm{Geo}(1/2)$ distributions): $\mathbb{E}\left(M_{n}\right)=\log(n)+O(1)$ and $\mathrm{Var}(M_{n})=\log(n)+O(1)$ and thus, similarly to the Morris Counter, there are only $\log(\log(n))+O(1)$ bits required on average to save the MaxGeo Counter after $n$ increment requests.
@@ 4.3 General Probabilistic Counting with Stochastic Averaging
Here we recall briefly a General Probabilistic Counting with Stochastic Averaging algorithm, based on the original idea from [23]. Assume that there are $m$, initially empty lots related to some independent copies of some probabilistic counter. For each increment request, we connect it to one of the groups uniformly at random. Finally, we perform incrementation requests separately and independently for each lot, obtaining the following. $M[1],M[2],\ldots,M[m]$.
@@
Then (according to [23]), for any $m=2^{k}$, $k\in\mathbb{N}$, $\mathbb{E}\left(\Xi_{n}(m)\right)\approx n\left(1+\frac{0.31}{m}\right)$ and $\mathrm{Var}(\Xi_{n}(m))=n^{2}\left(\frac{0.61}{m}\right).$
@@ 4.4 LogLog counter
In [13] a LogLog algorithm was proposed. It is based on $m=2^{k}$ counters $(M[j])_{j=1}^{m}$, where $k>0$. We may interpret this algorithm in the context of probabilistic counters. In such a scenario, it takes a hashed value (binary sequence) as input on every increment request. The first $k$ bits of the hash determine which of the $m$ counters should be incremented (the index $j$ is chosen as the decimal representation of the sequence restricted to these first $k$ bits; hence it translates the increment request of the LogLog counter to the increment request of one of the $m$ internal counters). Consider the first non-zero bit of a tail of the sequence (starting from $(k+1)$-th bit). Its position $R$ in this tail follows the $\mathrm{Geo}\left(\frac{1}{2}\right)$ distribution assuming the uniform distribution of the input sequences. If $R>M[j]$, then $M[j]$ should become $R$. Otherwise, it does not change.
@@
We are interested in the expectation and accuracy of the cardinality estimator, which can be briefly described as follows: $\mathbb{E}\left(\mathrm{LogLog}_{n}^{(m)}\right)\approx n$ and $\mathrm{Var}(\mathrm{LogLog}_{n}^{(m)})\approx\frac{1.69n^{2}}{m}$ (see [13] for more details). However, in order to control the LogLog counter, on average about $m\log\log\left(\frac{n}{m}\right)$ bits of memory are needed.
@@ 4.5 HyperLogLog
The maximum of geometric variables is also used as a primitive in the well known HyperLogLog algorithm (see [22]). Therefore, its privacy properties are important both from the theoretical and practical point of view. Essentially, in HyperLogLog we perform the general PCSA algorithm, but the final estimation is somehow different:
@@
where $\alpha_{k}$ is a constant dependent only on $k$ (see [22] for more details). It should be noted that HyperLogLog related algorithms (mentioned in Section 1) are the best-known procedures designated for cardinality estimation, and are close to optimum [30]. According to [22], for $m=2^{k}$, where $k\geqslant 4$,
@@ 5 Privacy-Preserving Survey via Probabilistic Counters
In this section, we present an example scenario for data aggregation using probabilistic counters. We assume that there is a server (alternatively, we call it aggregator) and a collection of nodes (e.g., mobile phone users), and we want to perform a boolean survey with a sensitive question. That is, each user sends $^{\prime}0^{\prime}$ if his answer is no and $^{\prime}1^{\prime}$ if the answer is yes. We assume that the connections between users and the server are perfectly secure, and the data can safely get to the trusted server. This can be performed using standard cryptographic solutions. The server’s goal is to publish the sum of all responses $^{\prime}1^{\prime}$ in a way that preserves privacy. This goal could obviously be achieved by simply collecting all the data and adding an appropriately calibrated Laplace noise (see [18]). However, we aim to show that probabilistic counters have inherently sufficient randomness to be differentially private without any auxiliary randomising mechanism.
@@
Note that we assume that the Adversary has access only to the released value. We also released only the counter value itself, which does not estimate the responses of $^{\prime}1^{\prime}$. This estimation is a function of the released value, which is different for Morris or MaxGeo Counter. There can also be various ways to estimate the actual number using a counter value. However, this does not matter for our case, as differential privacy is conveniently fully resilient to post-processing (see [18]). The graphical representation of our scenario considered is presented in Figure 2.
@@ 6.2 General result on Morris’ Counter privacy
Let $M$ denote the Morris Counter. If $c>0$ satisfies $\lceil c\log(\ln(n))\rceil\geqslant 1$, then $M$ is $\left(\varepsilon(n),\delta(n)\right)$-DP with parameters $\varepsilon(n)=O\left(\frac{\left(\log(n)\right)^{2}}{n}\right)$ and
$\delta(n)=O\left(n^{-\left(\ln(n)\right)^{c-1}}+n^{-1}\left(\ln(n)\right)^{-c}\right)$.
@@ 7.1 Discussion on arragements of the survey scenario
Consider a result of Morris Counter with a small number $n$ of increment requests (for example, a number of respondents suffering from a rare illness). Therefore, we will likely require the parameter $\varepsilon$ to be at most some threshold, e.g. $1$. Therefore, from Theorem 1, we should add $x$ counts where $L(x)\leqslant{16}/{(x-8)}\leqslant 1$, so $x\geqslant 24$. Note that we do not include $n$ in the above formula since it is not known in advance. Therefore, using the Morris Counter, the above survey is at least $(L(n+24),0.00033)$-DP. However, the estimator should be modified as well, i.e., $M^{\prime}_{n}=M_{n+24}$, so $\hat{n^{\prime}}=\max\{2^{M^{\prime}_{n}}-26,0\}$ (since $2^{M^{\prime}_{n}}$ may be smaller than $26$). margin:
On the other hand, using MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$ (see Theorem 3).
@@
Assume that we have at least $n=200$ increment requests. From Theorem 1, we have $L(n)\leqslant{16}/{(n-8)}\leqslant 0.08334$. Therefore, using the Morris Counter, the above survey is $(0.08334,0.00033)$-DP.
On the other hand, using the MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$. Here we present an example.
@@
Note that from a differential privacy perspective, both the general PCSA algorithm and HyperLogLog can be seen as arbitrary postprocessing performed on the $m$ MaxGeo counters. Moreover, since each response goes to one counter only, they are independent of each other, so we can use the parallel composition theorem (see [18]).
@@ 7.2 Comparison of Morris and MaxGeo Counters
Imagine that 100 million people participate in a general health survey with $100$ yes/no sensitive questions.
For every question, we would like to estimate the number of people who answered yes, but we want to guarantee the differential privacy property at a reasonable level. Realise that if the number of yes answers is very small for some questions (e.g., when the question is about a very rare disease), then the number of no answers may be counted instead. Obviously, this method gains privacy, but loses precision.
According to Figure 4, if we use the Laplace method, then we may need approximately $100\log(10^{8})=2657.54\ldots$ bits to store the counters. However, if we use the Morris Counter instead, about $100\log(\log(10^{8}))=473.20\ldots$ bits are needed. Note that all terms $O(1)$ in the ”Average memory” row of Figure 4 are bounded by $1$. Hence, its impact is negligible from a practical point of view.
One may also complain about the heavy use of the pseudo random number generator (PRNG) that probabilistic counters make. However, this problem may be resolved by generating the number of increment requests which have to be forgotten until the next update of the counter by using appropriate geometric distributions (see, for example, [33] for a similar approach applied to the reservoir sampling algorithm). This way, the use of PRNG can be substantially reduced.

### v7

@@ 3 Differential Privacy Preliminaries
In this section, we briefly recall differential privacy. For more details, see, e.g. Dwork and Roth (2014). We denote the set of (positive) natural numbers by $\mathbb{N}$ and the set of all integers by $\mathbb{Z}$. Moreover, let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $a,b\in\mathbb{Z}$ let us define a discrete interval $a,b]\cap\mathbb{Z}$ by $[a:b]$. We also define $n]=\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$. We assume that there exists a trusted curator who holds, or securely obtains, the data of individuals in a (possibly distributed) database $x$. Every row of $x$ consists of the data of some individual. By $\mathcal{X}$, we denote the space of all possible rows. The goal is to protect the data of every single individual, even if all users except one collude with an adversary to breach the privacy of this single, uncorrupted user. On the other hand, the curator is responsible for producing a release – a possibly accurate response to a requested query. This response is then released to the public, who is allowed to perform a statistical analysis on it. The differential privacy is, by design, resilient to post-processing attacks, so even if the adversary obtains the public release, he will not be able to infer anything about specific individuals participating in this release.
@@
Definition 2 (Differential Privacy – from Dwork and Roth (2014))
@@
Note that $\mathbb{E}\left(X\right)=0$ and $\mathrm{Var}(X)=2\lambda^{2}$.
Let $c(x)$ be the number of rows in $x$, which satisfy a given property. Note that $c$ in the differential privacy literature is usually referred to as count query. Imagine that an aggregating mechanism is defined as follows: $\mathcal{M}(x)=c(x)+\mathcal{L}(\varepsilon^{-1})$. Then $\mathcal{M}$ is $(\varepsilon)$-DP (for more precise properties of Laplace noise, see Dwork and Roth (2014)).
In the privacy analysis of large-scale distributed protocols, two types of approaches are typically distinguished: event/record-level (e.g., Wang et al. (2019); Papernot et al. (2017)), where information about a specific event/record is protected, and user-level privacy (e.g., McMahan et al. (2018)), which protects the privacy of the users themselves. Generally, the latter is a stronger model (see the discussion in McMahan et al. (2018)).
@@ 4 Probabilistic Counters — preliminaries
In Figure 1, one can see a graphical representation of the probabilistic counter. As mentioned, increment requests are indicated by $^{\prime}1^{\prime}$ and other rows by $^{\prime}0^{\prime}$ input. The dice represent randomness. The $\mathbf{X}$-mark indicates that there is no action.
@@ 4.1 Morris Counter
We begin with a short description of the Morris Counter (originally referred to as an approximate counter Morris (1978); Flajolet (1985)). Fix $a>1$. The algorithm 1 is a very simple pseudocode of the Morris Counter Morris (1978).
@@ 5    $M\leftarrow M+1$;
Roughly speaking, we start with $M=1$. Each incoming increment request triggers a random event. This event increments the counter ($M\leftarrow M+1$) with probability $a^{-M}$ ($r\sim\mathrm{Uni}([0,1])$ generates a number uniformly at random from the interval $[0,1]$, using in practice some Pseudo Random Number Generator). Note that this approximate counting protocol can be easily distributed. Indeed, any entity who wants to increment the counter only has to send the request to increment it. These requests can be queued on the server and resolved one after another. A detailed description of the approximate counting method can be found in Morris (1978); Flajolet (1985). Throughout this article, we examine only a standard Morris Counter i.e., with the base $a=2$. Morris Counter can also be defined recursively.
@@
The accuracy of Morris Counter has been thoroughly analysed in various classical papers. The first detailed analysis was proposed by Ph. Flajolet in Flajolet (1985). In this part, we present the essence of theorems presented in this paper, which will be useful later on.
@@ 4 $C\leftarrow\max\{C,r\}$;
The expectation and variance of the maximum of $n$ i.i.d. geometric variables have already been analysed in the literature. For instance, Szpankowski and Rego Szpankowski and Rego (1990) provided exact formulas for the’ expected value and variance of such variables. However, they are impractical for large applications $n$. Hence, they also provided asymptotics (here, for a maximum of $n$ independent $\mathrm{Geo}(1/2)$ distributions): $\mathbb{E}\left(M_{n}\right)=\log(n)+O(1)$ and $\mathrm{Var}(M_{n})=\log(n)+O(1)$ and thus, similarly to the Morris Counter, there are only $\log(\log(n))+O(1)$ bits required on average to save the MaxGeo Counter after $n$ increment requests.
@@ 4.3 General Probabilistic Counting with Stochastic Averaging
Here we recall briefly a General Probabilistic Counting with Stochastic Averaging algorithm, based on the original idea from Flajolet and Martin (1985). Assume that there are $m$, initially empty lots related to some independent copies of some probabilistic counter. For each increment request, we connect it to one of the groups uniformly at random. Finally, we perform incrementation requests separately and independently for each lot, obtaining the following. $M[1],M[2],\ldots,M[m]$.
@@
Then (according to Flajolet and Martin (1985)), for any $m=2^{k}$, $k\in\mathbb{N}$, $\mathbb{E}\left(\Xi_{n}(m)\right)\approx n\left(1+\frac{0.31}{m}\right)$ and $\mathrm{Var}(\Xi_{n}(m))=n^{2}\left(\frac{0.61}{m}\right).$
@@ 4.4 LogLog counter
In Durand and Flajolet (2003) a LogLog algorithm was proposed. It is based on $m=2^{k}$ counters $(M[j])_{j=1}^{m}$, where $k>0$. We may interpret this algorithm in the context of probabilistic counters. In such a scenario, it takes a hashed value (binary sequence) as input on every increment request. The first $k$ bits of the hash determine which of the $m$ counters should be incremented (the index $j$ is chosen as the decimal representation of the sequence restricted to these first $k$ bits; hence it translates the increment request of the LogLog counter to the increment request of one of the $m$ internal counters). Consider the first non-zero bit of a tail of the sequence (starting from $(k+1)$-th bit). Its position $R$ in this tail follows the $\mathrm{Geo}\left(\frac{1}{2}\right)$ distribution assuming the uniform distribution of the input sequences. If $R>M[j]$, then $M[j]$ should become $R$. Otherwise, it does not change.
@@
We are interested in the expectation and accuracy of the cardinality estimator, which can be briefly described as follows: $\mathbb{E}\left(\mathrm{LogLog}_{n}^{(m)}\right)\approx n$ and $\mathrm{Var}(\mathrm{LogLog}_{n}^{(m)})\approx\frac{1.69n^{2}}{m}$ (see Durand and Flajolet (2003) for more details). However, in order to control the LogLog counter, on average about $m\log\log\left(\frac{n}{m}\right)$ bits of memory are needed.
@@ 4.5 HyperLogLog
The maximum of geometric variables is also used as a primitive in the well known HyperLogLog algorithm (see Flajolet et al. (2007)). Therefore, its privacy properties are important both from the theoretical and practical point of view. Essentially, in HyperLogLog we perform the general PCSA algorithm, but the final estimation is somehow different:
@@
where $\alpha_{k}$ is a constant dependent only on $k$ (see Flajolet et al. (2007) for more details). It should be noted that HyperLogLog related algorithms (mentioned in Section 1) are the best-known procedures designated for cardinality estimation, and are close to optimum Indyk and Woodruff (2003). According to Flajolet et al. (2007), for $m=2^{k}$, where $k\geqslant 4$,
@@ 5 Privacy-Preserving Survey via Probabilistic Counters
In this section, we present an example scenario for data aggregation using probabilistic counters. We assume that there is a server (alternatively, we call it aggregator) and a collection of nodes (e.g., mobile phone users), and we want to perform a boolean survey with a sensitive question. That is, each user sends $^{\prime}0^{\prime}$ if his answer is no and $^{\prime}1^{\prime}$ if the answer is yes. We assume that the connections between users and the server are perfectly secure, and the data can safely get to the trusted server. This can be performed using standard cryptographic solutions. The server’s goal is to publish the sum of all responses $^{\prime}1^{\prime}$ in a way that preserves privacy. This goal could obviously be achieved by simply collecting all the data and adding an appropriately calibrated Laplace noise (see Dwork and Roth (2014)). However, we aim to show that probabilistic counters have inherently sufficient randomness to be differentially private without any auxiliary randomising mechanism.
@@
Note that we assume that the Adversary has access only to the released value. We also released only the counter value itself, which does not estimate the responses of $^{\prime}1^{\prime}$. This estimation is a function of the released value, which is different for Morris or MaxGeo Counter. There can also be various ways to estimate the actual number using a counter value. However, this does not matter for our case, as differential privacy is conveniently fully resilient to post-processing (see Dwork and Roth (2014)). The graphical representation of our scenario considered is presented in Figure 2.
@@ 6.2 General result on Morris’ Counter privacy
Let $M$ denote the Morris Counter. If $c>0$ satisfies $\lceil c\log(\ln(n))\rceil\geqslant 1$, then $M$ is $\left(\varepsilon(n),\delta(n)\right)$-DP with parameters $\varepsilon(n)=O\left(\frac{\left(\log(n)\right)^{2}}{n}\right)$ and $\delta(n)=O\left(n^{-\left(\ln(n)\right)^{c-1}}+n^{-1}\left(\ln(n)\right)^{-c}\right)$.
@@ 7.1 Discussion on arragements of the survey scenario
Consider a result of Morris Counter with a small number $n$ of increment requests (for example, a number of respondents suffering from a rare illness). Therefore, we will likely require the parameter $\varepsilon$ to be at most some threshold, e.g. $1$. Therefore, from Theorem 1, we should add $x$ counts where $L(x)\leqslant{16}/{(x-8)}\leqslant 1$, so $x\geqslant 24$. Note that we do not include $n$ in the above formula since it is not known in advance. Therefore, using the Morris Counter, the above survey is at least $(L(n+24),0.00033)$-DP. However, the estimator should be modified as well, i.e., $M^{\prime}_{n}=M_{n+24}$, so $\hat{n^{\prime}}=\max\{2^{M^{\prime}_{n}}-26,0\}$ (since $2^{M^{\prime}_{n}}$ may be smaller than $26$).
On the other hand, using the MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$ (see Theorem 3).
@@
Assume that we have at least $n=200$ increment requests. From Theorem 1, we have $L(n)\leqslant{16}/{(n-8)}\leqslant 0.08334$. Hence, using the Morris Counter, the above survey is $(0.08334,0.00033)$-DP.
On the other hand, using the MaxGeo Counter for a given $\varepsilon$ and $\delta$ we get $\left(\varepsilon,\delta\right)$-DP as long as $n\geqslant\dfrac{\ln(\delta)}{\ln\left(1-2^{-l_{\varepsilon}}\right)}$, where $l_{\varepsilon}=\left\lceil\log\left(1+{1}/{\varepsilon}\right)\right\rceil$.
@@
Note that from a differential privacy perspective, both the general PCSA algorithm and HyperLogLog can be seen as arbitrary postprocessing performed on the $m$ MaxGeo counters. Moreover, since each response goes to one counter only, they are independent of each other, so we can use the parallel composition theorem (see Dwork and Roth (2014)).
@@ 7.2 Comparison of Morris and MaxGeo Counters
Imagine that 100 million people participate in a general health survey with $100$ yes/no sensitive questions.For every question, we would like to estimate the number of people who answered yes, but we want to guarantee the differential privacy property at a reasonable level. Realise that if the number of yes answers is very small for some questions (e.g., when the question is about a very rare disease), then the number of no answers may be counted instead. Obviously, this method gains privacy, but loses precision.
According to Figure 4, if we use the Laplace method, then we may need approximately $100\log(10^{8})=2657.54\ldots$ bits to store the counters. Let us note however, if we use the Morris Counter instead, about $100\log(\log(10^{8}))=473.20\ldots$ bits are needed. Note that all terms $O(1)$ in the ”Average memory” row of Figure 4 are bounded by $1$. Hence, its impact is negligible from a practical point of view.
One may also complain about the heavy use of the pseudo random number generator (PRNG) that probabilistic counters make. However, this problem may be resolved by generating the number of increment requests which have to be forgotten until the next update of the counter by using appropriate geometric distributions (see, for example, Li (1994) for a similar approach applied to the reservoir sampling algorithm). This way, the use of PRNG can be substantially reduced.